OpenStudy (anonymous):

Find a function f '(x, y) that is harmonic in the region in the right half plane that lies between the curves x^2 − y^2 = 1 and and x^2 − y^2 = 4 and takes the value 2 on the left curve and value 5 on the right.

OpenStudy (anonymous):

Any math wizzes around? i've come to find that Re( a Z^2+b) = a (x^2-y^2) +b assuming a,b are elements of R

OpenStudy (anonymous):

using a substitution we can find two equations given (1)..... a x 1 +b =2 and (2)....a x 4 +b =5